### v1

# From Seeing to Doing: Bridging Reasoning and Decision for Robotic Manipulation

Abstract
Achieving generalization in robotic manipulation remains a critical challenge, particularly for unseen scenarios and novel tasks. Current Vision-Language-Action (VLA) models, while building on top of general Vision-Language Models (VLMs), still fall short of achieving robust zero-shot performance due to the scarcity and heterogeneity prevalent in embodied datasets. To address these limitations, we propose FSD (From Seeing to Doing), a novel vision-language model that generates intermediate representations through spatial relationship reasoning, providing fine-grained guidance for robotic manipulation. Our approach combines a hierarchical data pipeline for training with a self-consistency mechanism that aligns spatial coordinates with visual signals. Through extensive experiments, we comprehensively validated FSD’s capabilities in both “seeing” and “doing”, achieving outstanding performance across 8 benchmarks for general spatial reasoning and embodied reference abilities, as well as on our proposed more challenging benchmark VABench. We also verified zero-shot capabilities in robot manipulation, demonstrating significant performance improvements over baseline methods in both SimplerEnv and real robot settings. Experimental results show that FSD achieves 54.1% success rate in SimplerEnv and 72% success rate across 8 real-world tasks, outperforming the strongest baseline by 30%. More visualizations are available on website.
1 Introduction
A driving force behind robotics research is the pursuit of generalization: creating agents capable of versatile action across diverse robotic platforms, extending beyond familiar tasks, objects, and environments while adapting to dynamic visual inputs. Current approaches [21, 3] leverage pre-trained Vision-Language Models (VLMs) and transform them into Vision-Language-Action Models (VLAs) using large-scale embodied datasets. This enables systems to interpret natural language instructions and generate robotic manipulation actions [2]. The intention is to capitalize on the generalization capabilities of VLMs pre-trained on internet-scale data, with the hope that resulting VLAs will adapt to novel scenarios involving unseen objects and tasks. However, empirical evidence [72, 67, 35] demonstrates that directly applying the generalization power of VLMs falls short of achieving strong zero-shot performance on completely novel tasks.
Figure 1: Overview of FSD. FSD unlocks visual aids reasoning and generation through Spatial Relationship-Focused CoT, demonstrating exceptional generalization capabilities that enable zero-shot robot manipulation and achieving remarkable performance across multiple benchmarks.
We attribute the limited generalization in VLA-based systems to scarcity and heterogeneity of datasets. Robotics data remains limited compared to language and vision datasets, preventing similar scaling laws [18, 29]. Despite growth in embodied datasets [40], their insufficient coverage and diversity prevent robust zero-shot generalization. Additionally, robotic embodiment heterogeneity [54] causes significant variation in VLA outputs, making end-to-end supervised learning from vision and language to diverse action outputs a potentially unrealistic path toward generalization.
We introduce a novel pipeline addressing generalization challenges in robotic manipulation. Our approach leverages VLMs’ visual understanding capabilities, augmented with step-by-step reasoning to extract unified mid-level structural representations independent of robot embodiment. This representation is key to generalizing learning across diverse physical interactions and dynamic behaviors. Specifically, this mid-level representation includes spatial affordance boxes/points and visual traces, each represented as marked coordinates within visual images. These visual aids provide expressive yet compact spatial information that enables effective reasoning and decisions, overcoming both scarcity and heterogeneity limitations. We introduce FSD (From Seeing to Doing), a model that generates visual intermediate representations through spatial understanding and reasoning (Fig. 1). The FSD model comprises three key components: (1) Spatial Relationship-Focused Visual Chain-of-Thought (SrCoT), which conducts multi-step reasoning anchored by object coordinates and spatial relationships, treating visual aid generation as a reasoning process; (2) A hierarchical data pipeline combining large-scale embodied datasets with common sense data, establishing a weak-to-strong capability enhancement training process; and (3) A self-consistency mechanism that aligns understanding and generation by binding spatial coordinates with specific visual signals. We employ supervised fine-tuning of VLMs using annotated spatial COT datasets for visual aids generation, then action executors follow spatial affordances and visual traces through simple planning methods to complete action execution.
Our FSD model generalizes effectively to new instructions, and scenes through enhanced reasoning abilities. Our contributions include: (1) A novel paradigm bridging VLM reasoning with embodied decisions via visual aids; (2) The SrCoT method enabling multi-step reasoning for visual aid generation and guiding zero-shot manipulation; (3) Our weak-to-strong spatial reasoning and visual aids datasets, along with VABench, a manually annotated challenging benchmark for visual aids generation; (4) Superior performance across 8 benchmarks in spatial reasoning, free space reference, and visual aids generation, with zero-shot deployment achieving 54.1% success in SimplerEnv and 73% in 8 real-world tasks, outperforming mainstream VLA models by 30%.
2 Related Work
Spatial Understanding and Reasoning with VLMs Spatial understanding and reasoning [32, 48, 8, 28, 44] require VLMs to infer spatial information beyond 2D RGB images, a capability crucial for embodied AI applications such as navigation [48, 12, 22] and manipulation [66]. Recent advances include SpatialVLM [5], which explicitly incorporates spatial primitives and coordinate systems to enhance geometric reasoning capabilities. Similarly, SpatialRGPT [7] and SpatialBot [4] improve spatial capabilities through more precise spatial relationship modeling. FSD specifically targets embodied manipulation scenarios by enhancing spatial reasoning capabilities through novel SrCoT mechanisms and sophisticated self-consistency alignment techniques.
Visual Chain-of-thought Reasoning The integration of Chain-of-thought (CoT) [56] and its variants [69, 63, 62] has significantly enhanced LLM reasoning abilities through structured step-by-step processes. For the multimodal reasoning challenge, researchers have developed various CoT approaches [37, 70, 61, 58] that establish appropriate reasoning anchors and extend reasoning pathways. Shikra [6] improves referential expression by incorporating specific visual regions into the reasoning process, while VoCoT [27] extends visually-grounded reasoning chains. VisualCoT [46] provides a comprehensive benchmark validating CoT effectiveness in multi-hop image reasoning tasks. EmbodiedCoT [67] pioneered CoT application in embodied AI by enhancing intermediate reasoning through fine-tuning OpenVLA [21]. In contrast, FSD uniquely integrates visual-spatial reasoning using spatial relationship graphs as reasoning anchors.
Visual Aids Empowered Robotic Manipulation Extracting embodiment-agnostic visual aids to enhance training efficiency has emerged as a promising paradigm in robotic manipulation. Numerous studies [1, 57, 59, 72, 65] have explored robotic policy learning based on visual traces, though most remain confined to specific tasks with cross-embodiment applicability. LLaRVA [39] advances this field by predicting visual traces to better align visual and action spaces for robot learning, compiling an impressive large visual trace dataset, yet struggles to generalize to novel downstream tasks without task-specific fine-tuning. Spatial affordance represents another effective visual aid, with several works [66, 38, 42, 48, 16, 60, 26] demonstrating its utility in robotic manipulation tasks. Our key insight is that the scarcity of embodied data fundamentally limits purely data-driven visual aid prediction approaches, impeding zero-shot generalization to unseen scenarios. Therefore, we employ a reasoning-driven paradigm to activate general spatial abilities and enhance generalization.
3 Bridging Reasoning and Decision through Visual Aids Generation
To harness VLMs’ visual perception capabilities for cross-domain generalization, we introduce FSD (From Seeing to Doing)—a model exhibiting robust spatial reasoning and generation abilities. We first establish visual aids through spatial affordances and visual traces, then develop Spatial Relationship-Focused CoT (SrCoT), which leverages object-centric coordinates and their spatial relationships as reasoning anchors. Supporting this approach requires precise spatial understanding and complex instruction-following capabilities to generate coordinate representations. We implement a progressive weak-to-strong data construction pipeline across five capability levels, complemented by a self-consistency alignment mechanism that enhances understanding and generation abilities.
3.1 Definition of Visual Aids
Figure 2: Diagrams of Visual Aid Types
As shown in Fig. 2, FSD utilizes three visual aids of increasing complexity, all defined within normalized image coordinates $\mathbf{x}=(p,q)\in[0,1000]^{2}\subset\mathbb{R}^{2}$: Spatial affordance boxes $\mathcal{B}=[x_{1},y_{1},x_{2},y_{2}]$ define target regions for object placement. For instructions like "Put the sushi into the silver pot," the model must infer coordinates for unmarked free space—beyond standard detector capabilities. Spatial affordance points $\mathcal{P}=\{(x_{i},y_{i})\mid i=1,2,...,n\}$ provide more precise and flexible placement with reduced redundancy. Object-centric visual traces $\boldsymbol{\tau}=\{\mathbf{x}_{t}\mid t=1,2,...,T\}$ represent ordered coordinate sequences that describe manipulation trajectories, where $T$ denotes the sequence length. These traces enable complex instruction execution, cross-embodiment transfer, and collision avoidance. We implement these representations in 2D rather than 3D due to limited high-quality 3D data availability [68]. Our object-centric rather than agent-centric visual aids approach effectively circumvents the limitations of heterogeneous embodied data while leveraging general visual datasets without robots, thus enabling robust generalization to novel scenarios and tasks.
Figure 3: Inspired by the process of human reasoning, FSD uses a spatial relationship graph as an anchor to derive a visual chain-of-thought reasoning process for visual trace generation.
3.2 Spatial Relationship-Focused Visual Chain-of-thought
To enable VLMs to generate spatial visual aids, a direct approach is supervised fine-tuning [39, 66, 23, 67] using these aids as a new action space or employing generative models [47, 59]. However, the heterogeneity and scarcity of embodied datasets limit this method. Directly aligning RGB images with coordinate points is challenging and prone to overfitting, hindering generalization. How can we stimulate VLMs’ spatial reasoning abilities to guide the generation rather than merely relying on extensive demonstrations? Inspired by human cognition (Fig. 3 (Top)), when executing tasks like "putting broccoli into a pot," humans first locate relevant objects, then plan movement paths based on relative positions while considering feasibility and obstacles. During this process, humans construct reasoning chains, repeatedly referencing object positions and establishing spatial relationships.
Based on these considerations, we introduce Spatial Relationship-Focused Visual Chain-of-thought (SrCoT). This approach guides VLMs to generate visual aids through structured reasoning based on spatial relationship graphs. SrCoT consists of two essential phases: 1 Description: We generate object-centric region captions establishing a task-relevant spatial relationship graph where nodes represent objects with their coordinates and edges denote relative relationships (above, below, left, right, etc.). 2 Reasoning: Using the spatial relationship graph as anchor points, we determine start and end coordinates through object references and free space reasoning, then iteratively derive intermediate points with explicit logical connections between steps. Thus, we prescribe a templated reasoning path for VLMs, enabling FSD to perform analogical reasoning in the spatial domain. While VLMs struggle to directly map future actions to image coordinates, our method leverages known object relationships as reference points for multi-hop analysis, simplifying the reasoning process. Fig. 3 (Bottom) demonstrates a complete reasoning sequence. This step-by-step SrCoT approach, though powerful, fundamentally depends on precise spatial understanding capabilities.
3.3 Weak-to-Strong Capability Dataset Construction Pipeline
The SrCoT mechanism demands enhanced capabilities from VLMs, including precise reference grounding, spatial understanding, and complex instruction-following capabilities (directly predicting future point trajectories) where mainstream VLMs [34, 30, 44, 8] show limitations. Consequently, we designed a weak-to-strong data construction pipeline to progressively develop these abilities.
FSD encompasses five hierarchical capability levels: 1 Region Grounding enables robots to focus on key objects in scenes. Although grounding capabilities have been broadly integrated into current VLMs [6, 64], understanding various small objects and complex scenes for embodied tasks is still limited; 2 Spatial Relationship understanding establishes prerequisite knowledge for spatial reasoning, forming the anchor points for SrCoT; 3 Spatial Reasoning builds upon these foundations to perform multi-hop analysis of object positions and relationships; and finally, 4 Spatial Affordance Generation and 5 Visual Trace Generation create actionable spatial guidance. Notably, SrCoT functions as a general visual-spatial reasoning mechanism applicable beyond visual traces to diverse spatial reasoning tasks. Through hierarchical spatial capability training, we enhance VLMs’ general spatial reasoning abilities, extending well beyond just embodied domains.
Figure 4: FSD screens data from large-scale embodied datasets, generates ground truth spatial relationship graph. We finally collected 300K data for 10+ embodiments with 5-level capabilities.
Automatic Dataset Construction: We leveraged extensive robot data from BridgeDataV2 [53], RT-X [40], and Droid [20] to construct FSD’s training dataset. Inspired by SpatialVLM [5] and SpatialRGPT [7], our automated data construction pipeline created 300K Supervised Fine-Tuning (SFT) data across numerous formats. After filtering demonstrations with unclear instructions, we used GPT4o [15] to nominate task-relevant objects while excluding out-of-range or overly complex items. We then built around objects using GroundedSam [45] for bounding boxes and segmentation masks (Level 1 dataset). For spatial relationship labels, we reconstructed 3D semantic scene graphs using Metric3Dv2 [13] for depth estimation, along with WildCamera [73] and PerspectiveFields [17] for camera parameters. This enabled 2D-3D mapping and spatial relationship graph construction. (Level 2 dataset). We only generate relative depth sorting data to infer positional relationships, so the accuracy requirement is not high. To improve data quality, in particular, we selected objects with a relative depth gap of 20% for subsequent generation. Afterward, we randomly sample spatial relationship graphs to construct spatial reasoning QA. We provide image captions, object coordinates, and relationships as a context for GPT4o to create complex QAs (Level 3 Dataset).
A core aspect of the FSD dataset is the visual aids generation. We employ a simple method with successful human demonstrations from embodied datasets and infer the process from the results. Spatial affordance represents the designated completion area for manipulation tasks. To create spatial affordance labels (Level 4 Dataset), we extract the manipulated object’s final position from the terminal frame, combine it with reference object positioning, calculate the precise affordance region, and re-render this information onto the initial frame. For visual trace generation (Level 5 Dataset), we employ a two-stage approach: first applying self-supervised keypoint extraction [14] to identify grasp points on manipulated objects, then utilizing Cotracker [19] to capture temporal dynamics from human demonstrations, subsequently projecting these trajectories onto the initial frame. Throughout this process, we employed strict rule-based filters and continually validated our approach against manually annotated test sets, iteratively refining our filtering criteria based on empirical feedback to ensure the resulting dataset met our quality requirements. The dataset presentation, data filtering process, and prompts used to generate the data are provided in Appendix A.
3.4 Self-Consistent Alignment for Spatial Understanding and Generation
High-quality SFT datasets enable VLMs to generate visual aids [66], yet these models struggle to understand the physical meaning of such annotations since coordinate spaces never appeared in pretraining data. The alignment between image coordinates and actual spatial positions presents a significant challenge. Therefore, we propose a self-consistency mechanism to further align FSD capabilities in spatial understanding and generation. We frame generation tasks inversely as understanding problems: if the forward task requires inferring visual trace $\boldsymbol{\tau}$ from an image $X_{v}$ and task instruction $X_{q}$, i.e. $(X_{v},X_{q})\rightarrow\boldsymbol{\tau}$, we construct the inverse task of predicting possible instructions given an image and visual traces $(X_{v},\boldsymbol{\tau})\rightarrow X_{q}$. This bidirectional approach helps the model comprehend spatial coordinates’ meanings and aligns coordinate space with image-text modalities, unifying visual aids as both understanding and generation signals while enhancing FSD spatial reasoning capabilities.
4 Training and Action Execution of FSD
Training: We follow the instruction tuning pipeline proposed by LLaVA-1.5 [33, 34]. As shown in Fig. 1, FSD’s architecture comprises an image encoder (CLIP-ViT-L-336px [11]), a linear projector, a language tokenizer, and a LLM (Vicuna-13B [71]). The image encoder processes images into tokens. These visual tokens are then projected into the same embedding space as language tokens through a two-layer linear projector. Only the projector and LLM weights are updated during fine-tuning while the vision encoder and tokenizer remain frozen. We built upon ASMv2 [55] as our foundation, which already incorporates basic relation conversation and reference grounding capabilities. The training process of FSD is divided into two stages: General Spatial Reasoning Enhancement: Using data from levels 1-3, we focus on improving the model’s embodied spatial reasoning capabilities. Following Yuan et al. [66] and Brohan et al. [3], we discovered that an appropriate data mixture is crucial for downstream performance. Joint training with mixed robotic and internet data ensures the model retains knowledge acquired during pre-training. Consequently, our instruction tuning utilizes a diverse 1.4M sample mixture including general visual question answering (VQA) data. This comprehensive training ensures FSD maintains robust general spatial knowledge while developing embodied capabilities. Visual Aids Generation and Understanding: Using data from levels 4-5 with the self-consistency mechanism, we specifically train visual aids generation and understanding abilities. FSD predicts a fixed set of 8 points for simplification when generating spatial visual traces. Additional training details and the summary of mixture datasets are provided in Appendix B.
Action Execution: FSD can reason from initial or intermediate task steps, freely selecting needed visual aids. When using bounding boxes, we sample the center as the target point; with affordance points, we directly sample one point. For visual trace execution, we first generate 2D visual traces $\boldsymbol{\tau}$ and obtain preliminary depth information from depth cameras. Following the pinhole camera model, we employ depth-based back-projection to map these to 3D space, yielding $\boldsymbol{\tau}^{3d}=\{\mathbf{x}_{t}^{3d}\mid t=1,2,...,T\}$. Next, based on the spatial position of the first point $\mathbf{x}_{1}$, we query GraspNet’s [9] grasp candidates $G$ to match the nearest grasp pose $G^{*}$. For relatively fixed scenes, we may also use predetermined grasp poses. Subsequently, we optimize the path trajectory using gradient descent-based interpolation, generating complete motion trajectories in SE(3) space, enabling the robotic arm to follow the 3D visual trajectory. When using only spatial affordance, we utilize CuRobo [49] as the motion planner to determine execution trajectories $T$ based on the target position $\mathcal{P}$. More details are provided in Appendix C. Unlike methods such as LLARVA [39] and EmbodiedCOT [67] which also utilize visual auxiliary aids, FSD transforms prediction tasks into reasoning tasks, better leveraging visual-spatial common knowledge without requiring scenario-specific fine-tuning.
5 Visual Aids Generation Benchmark
Few datasets exist for evaluating visual aid generation, Where2Place [66] provides 100 real-world images from homes and offices but is limited to direct and simple language instructions. And no benchmarks for trajectory prediction. To address this gap, we propose the Visual Aids Generation Benchmark (VABench). we manually annotated 300 problems from real-world and simulation datasets (OXE [40], BridgeData [53], Droid [20], Libero [31]), requiring models to infer visual aids given only natural language instructions similar to everyday human commands. For spatial affordance (VABench-Point), we measure the proportion of points falling within target regions. For models that only output bounding boxes, we sample uniformly within boxes. For visual trace (VABench-VisualTrace), we compute MAE and RMSE between ground truth $\boldsymbol{\tau}=\{\mathbf{x}_{t}\mid t=1,2,...,T\}$ and predictions $\hat{\boldsymbol{\tau}}=\{\hat{\mathbf{x}}_{t}\mid t=1,2,...,\hat{T}\}$: $\text{MAE}=\frac{1}{T}\sum_{t=1}^{T}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|$, $\text{RMSE}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}}$. We interpolate when trajectory lengths differ and normalize coordinates to a 1000×1000 space for consistent evaluation. Since multiple valid solutions exist for each instruction, we simulated human evaluation standards by establishing detailed assessment criteria and added metrics with MLLM-based qualitative scoring (0-10) of visualized trajectories, named GPT Score. We provide detailed dataset information and evaluation procedures in Appendix D.
Figure 5: FSD directly generates visual aids based on task instructions for novel tasks and scenarios. 1st row: affordance bounding boxes; 2nd row: affordance points; 3rd and 4th rows: visual traces.
6 Experiments
We evaluated FSD across two dimensions: Seeing and Doing. For Seeing, we tested its general spatial reasoning and visual aids generation capabilities. For Doing, we conducted zero-shot manipulation experiments in both SimplerEnv [25] simulation and real-world xArm robotic platforms to assess its practical generalization performance.
6.1 Evaluation of Spatial Understanding and Reasoning Capabilities
Benchmarks and Baselines. General Spatial Reasoning Capabilities: We evaluated general spatial reasoning capabilities using five popular benchmarks: CVBench [51], BLINK [10], CRPE [55], SAT [44], and EmbSpatial-Bench [8]. These benchmarks encompass 15 subtasks measuring various spatial competencies including counting, spatial relationships, distance estimation and so on. We included two leading closed-source models: GPT-4o [15] and GPT-4V as performance reference. Subsequently, we conducted comparative analyses against other open-source spatial enhanced MLLMs, including LLaVA-1.5 [33], SAT-Dynamic [44], RoboPoint [66], and ASMv2 [55], all with 13B parameters. Object and Free Region Reference Capabilities: Following Yuan et al. [66], we assessed embodied spatial capabilities using the RoboRefIt [36] and Where2Place [66] benchmarks. We compared mainstream closed-source models and MLLM for enhancing spatial abilities (SpatialBot [4], SpaceLLaVA [5], and RoboBrain [16]). We used the proportion of predicted points within specified objects/regions as the accuracy metric. For models without point output support, we asked models to output bounding boxes of target objects/free space regions, then sampled evenly within these bounding boxes. Spatial Affordance and Visual Trace Capabilities: We utilized our VABench to evaluate the capabilities. We found very few models with this capability for Visual Trace prediction, so we trained an end-to-end prediction baseline model using a pre-trained DINOv2 [41] encoder coupled with a multi-layer transformer [52] architecture to predict visual trajectories, trained on the same visual trajectory data, named DINOv2 Predictor. We conducted this comparison to demonstrate the advantages of our reasoning-based FSD approach. More detailed descriptions of benchmarks and baselines are provided in Appendix E.
FSD exhibits superior general spatial reasoning capabilities. As shown in Table 1, FSD achieves a top ranking of 1.3 across 18 subtasks from 5 spatial benchmarks, surpassing other 13B VLMs and competing with GPT-4o. FSD demonstrates particular strengths in 3D depth perception (88.0%), 3D distance estimation (86.7%), and spatial relationship (78.3%). This confirms the effectiveness of our data mixing strategy and capability construction in enhancing spatial reasoning abilities. We believe stronger spatial foundational capabilities enable enhanced embodied perception.
FSD excels in object reference and free space localization. The results in Table 2 demonstrate FSD’s ability to accurately identify objects and free spaces from language instructions. For object reference (RoboRefIt), FSD achieves 56.7% accuracy, surpassing both GPT-4o (15.3%) and specialist models like RoboPoint (49.8%) by significant margins. On the more challenging free space reference task (Where2Place), FSD performs competitively with RoboPoint (45.8% vs. 46.0%) while substantially outperforming other models. This improvement stems from enhanced spatial understanding through our SrCoT mechanism. See Appendix F for more visualizations.
Table 1: Performance comparison on 5 spatial reasoning benchmarks. Bold and underlined values show best and second-best performance among open-source models.
CVBench CRPE SAT BLINK EmbSpatial Rank
Count 2DRel 3DDep 3DDis Avg. Exist. Subj. Pred. Obj. Avg. Val Real Count MV RelDepth SpRel Avg. Test
Closed-source models
GPT-4V 62.4 71.1 79.8 68.3 70.4 90.6 76.7 65.1 68.5 75.2 44.8 50.7 60.8 55.6 59.7 72.7 62.2 36.1 -
GPT-4o 65.9 85.5 87.8 78.2 79.4 93.3 81.9 71.8 73.6 80.2 49.4 57.5 49.2 60.2 74.2 69.2 63.2 49.1 -
Open-source models
LLaVA-1.5-13B 58.2 46.6 53.0 47.8 51.4 88.7 57.4 54.2 55.2 63.9 51.4 41.6 45.0 41.4 53.2 69.9 52.4 35.1 4.8
SAT-Dynamic-13B 61.5 89.7 80.7 73.0 76.2 87.5 60.6 57.6 65.2 67.7 87.7 54.9 35.8 44.4 73.4 66.4 55.0 51.3 2.8
RoboPoint-13B 56.5 77.2 81.5 57.7 68.2 93.2 66.3 62.4 70.9 73.2 53.3 46.6 48.3 44.4 62.1 65.7 55.1 51.4 2.8
ASMv2-13B 58.9 68.9 68.9 68.9 66.4 92.1 69.2 59.0 65.3 71.4 63.9 46.7 59.2 44.4 56.5 65.0 56.3 57.4 3.1
FSD-13B 62.4 86.5 88.0 86.7 80.9 94.0 75.2 65.1 70.4 76.2 73.2 63.3 60.0 46.6 70.2 78.3 63.8 63.3 1.3
Table 2: Performance comparison on object/free space reference benchmarks. The best results are highlighted in bold.
Benchmark GPT-4o SpaceLLaVA LLaVA-NeXT-34B SpatialBot-3B ASMv2-13B RoboBrain-7B RoboPoint-13B FSD-13B
RoboRefIt 15.3 21.3 19.9 23.6 48.4 10.1 49.8 56.7
Where2Place 29.1 11.8 15.0 15.0 22.0 16.6 46.0 45.8
Figure 6: Example of the FSD Reasoning Process FSD performs point-by-point reasoning and localization, ultimately generating a feasible visual trace.
Figure 7: Real-world robotic manipulation tasks performed by an xArm 6 robot (left) and performance comparison of FSD against baseline models GPT-4o and RoboPoint (right).
Table 3: Performance comparison on VABench. The best results are highlighted in bold.
(a) VABench-Point
Model Accuracy $\uparrow$
GPT4o 9.30
ASMv2 10.07
RoboPoint 19.09
FSD 61.82
   w/o SrCoT 26.21
   w/o Alignment 55.92
(b) VABench-VisualTrace
[…]
DINOv2 Predictor 128.32 117.49 2.17
FSD 78.26 63.44 5.49
   w/o SrCoT 99.53 80.06 4.00
   w/o Alignment 80.48 66.80 5.45
Table 4: SimplerEnv Evaluation on WidowX Robot The results of baselines are derived from [43]. ZS: zero-shot, FT: fine-tuning using BridgeData. Each task is tested 24 episodes.
Model Spoon→Towel Carrot→Plate Green→Yellow Eggplant→Basket Avg.
Grasp Succ. Grasp Succ. Grasp Succ. Grasp Succ.
RT-1-X [40] 16.7% 0% 20.8% 4.2% 8.3% 0% 0.0% 0% 1.1%
Octo-S [50] 77.8% 47.2% 27.8% 9.7% 40.3% 4.2% 87.5% 56.9% 30.0%
OpenVLA [21] 4.1% 0% 33.3% 0% 12.5% 0% 8.3% 4.1% 1.0%
RoboVLM (ZS) [24] 37.5% 20.8% 33.3% 25.0% 8.3% 8.3% 0.0% 0% 13.5%
RoboVLM (FT) [24] 54.2% 29.2% 25.0% 25.0% 45.8% 12.5% 58.3% 58.3% 31.3%
SpatialVLA (ZS) [43] 25.0% 20.8% 41.7% 20.8% 58.3% 25.0% 79.2% 70.8% 34.4%
SpatialVLA (FT) [43] 20.8% 16.7% 29.2% 25.0% 62.5% 29.2% 100% 100% 42.7%
FSD 58.3% 41.6% 58.3% 50.0% 91.6% 33.3% 37.5% 37.5% 54.1%
FSD demonstrates breakthrough performance in visual aid generation. As shown in Table Table 3(b), FSD significantly outperforms all baseline methods in generating precise spatial affordances and visual traces. For VABench-Point, FSD achieves 61.82% accuracy—three times higher than RoboPoint (19.09%), while for VABench-VisualTrace, FSD demonstrates significantly lower error rates and better GPT Score compared to the DINOv2 Predictor. Ablation studies confirm the critical contributions of both SrCoT and self-consistency alignment, validating our reasoning-based approach where step-by-step spatial analysis enables more accurate predictions than purely data-driven methods. We provide detailed comparisons with RoboBrain [16] in Appendix I.
Visualization of visual aids generated by FSD We present examples of generation in Fig. 5. FSD is able to adapt to diverse scenes, perspectives, and tasks, and the three types of visual traces generated can well match the task instructions. We also presented the reasoning process of FSD in Fig. 6, where the CoT provides clear guidance for the results. More visualizations are in Appendix F.
6.2 Evaluate the Decision Capability
SimplerEnv Simulation SimplerEnv [25] is a simulation environment specifically designed for evaluating real-world robotic manipulation. It provides a standardized platform, emphasizing reproducibility with real-world scenarios. We utilized FSD to generate visual traces and perform zero-shot deployment on the WidowX robotic arm. We compared its performance against state-of-the-art general manipulation VLA, including Octo [50], OpenVLA [21], RoboVLM [24] and SpatialVLA [43]. The results are presented in Table 4. FSD surpasses the baseline in most tasks without FT with a 54.1% success rate, demonstrating the strong generalization ability of FSD. Notably, FSD achieves the best average performance, even outperforming VLA models that even outperforming specially fine-tuned VLA models. This superior generalization capability demonstrates that visual aids representations provide better zero-shot transfer compared to direct action prediction approaches, bridging the gap between perception and control more effectively.
Real World Robot Evaluation As shown in Fig. 7, we conducted zero-shot tests with FSD on an xArm 6 robot across 8 tabletop manipulation tasks. The setup included an Intel RealSense L515 LiDAR camera. To test the capabilities of different visual aids, we used visual trace for the sponge and folding tasks, while affordance points were used for other tasks. We compared against RoboPoint and GPT-4o baselines, which were limited to predicting only start/end points rather than full trajectories. Robopoint often makes incorrect predictions in tasks involving spatial understanding. Under zero-shot conditions, FSD achieved 72% success rate, outperforming the strongest baseline by more than 30%. Notably, FSD successfully completed complex tasks with visual trace generation, e.g. cloth folding, which was beyond baseline capabilities. Full results are presented in Fig. 7 and we refer to Appendix G for detailed setup and more visualizations.
7 Conclusion
In this paper, we introduced FSD (From Seeing to Doing), bridging visual reasoning and robotic manipulation through intermediate spatial representations. Our approach addresses the challenges of data scarcity and embodiment heterogeneity through three key innovations: Spatial Relationship-Focused Visual Chain-of-Thought for multi-step reasoning; A hierarchical weak-to-strong data pipeline; and A self-consistency mechanism aligning spatial coordinates with visual signals. Experiments demonstrate FSD’s superior performance across multiple spatial reasoning visual aid generation benchmarks. In zero-shot robotic deployment, FSD achieved 72% success rate across diverse tasks, outperforming baselines by 30%. Limitations include reliance on 2D rather than 3D trajectory generation and constraints from available training data quality. Future works are detailed in Appendix J.

### v2

@@
 To address these limitations, we propose FSD (From Seeing to Doing), a novel vision-language model that generates intermediate representations through spatial relationship reasoning, providing fine-grained guidance for robotic manipulation.
-Our approach combines a hierarchical data pipeline for training with a self-consistency mechanism that aligns spatial coordinates with visual signals.
+Our approach combines a hierarchical data construction pipeline for training with a self-consistency mechanism that aligns spatial coordinates with visual signals.
 Through extensive experiments, we comprehensively validated FSD’s capabilities in both “seeing” and “doing”, achieving outstanding performance across 8 benchmarks for general spatial reasoning and embodied reference abilities, as well as on our proposed more challenging benchmark VABench.
 We also verified zero-shot capabilities in robot manipulation, demonstrating significant performance improvements over baseline methods in both SimplerEnv and real robot settings.
-Experimental results show that FSD achieves 54.1% success rate in SimplerEnv and 72% success rate across 8 real-world tasks, outperforming the strongest baseline by 30%.
-More visualizations are available on website.
+Experimental results show that FSD achieves 40.6% success rate in SimplerEnv and 72% success rate across 8 real-world tasks, outperforming the strongest baseline by 30%.
+More visualizations and datasets are available on website.
+Keywords: General Robotic Manipulation, VLM Reasoning, Spatial Chain-of-Thought, Spatial VLMs
+= Projects: https://embodied-fsd.github.io/
+[Uncaptioned image] Code Repository: https://github.com/pickxiguapi/Embodied-FSD
+= Datasets: https://huggingface.co/IffYuan
+= Contact: yuanyf@tju.edu.cn
 1 Introduction
 A driving force behind robotics research is the pursuit of generalization: creating agents capable of versatile action across diverse robotic platforms, extending beyond familiar tasks, objects, and environments while adapting to dynamic visual inputs.
-Current approaches [21, 3] leverage pre-trained Vision-Language Models (VLMs) and transform them into Vision-Language-Action Models (VLAs) using large-scale embodied datasets.
-This enables systems to interpret natural language instructions and generate robotic manipulation actions [2].
+Current approaches Kim et al.
+(2024), Brohan et al.
+(2023) leverage pre-trained Vision-Language Models (VLMs) and transform them into Vision-Language-Action Models (VLAs) using large-scale embodied datasets.
+This enables systems to interpret natural language instructions and generate robotic manipulation actions Black et al.
+(2024).
 The intention is to capitalize on the generalization capabilities of VLMs pre-trained on internet-scale data, with the hope that resulting VLAs will adapt to novel scenarios involving unseen objects and tasks.
-However, empirical evidence [72, 67, 35] demonstrates that directly applying the generalization power of VLMs falls short of achieving strong zero-shot performance on completely novel tasks.
+However, empirical evidence Zheng et al.
+(2024), Zawalski et al.
+(2024), Liu et al.
+(2024c) demonstrates that directly applying the generalization power of VLMs falls short of achieving strong zero-shot performance on completely novel tasks.
 Figure 1: Overview of FSD.
@@
 We attribute the limited generalization in VLA-based systems to scarcity and heterogeneity of datasets.
-Robotics data remains limited compared to language and vision datasets, preventing similar scaling laws [18, 29].
-Despite growth in embodied datasets [40], their insufficient coverage and diversity prevent robust zero-shot generalization.
-Additionally, robotic embodiment heterogeneity [54] causes significant variation in VLA outputs, making end-to-end supervised learning from vision and language to diverse action outputs a potentially unrealistic path toward generalization.
-We introduce a novel pipeline addressing generalization challenges in robotic manipulation.
+Robotics data remains limited compared to language and vision datasets, preventing similar scaling laws Kaplan et al.
+(2020), Lin et al.
+(2024a).
+Despite growth in embodied datasets O’Neill et al.
+(2023), their insufficient coverage and diversity prevent robust zero-shot generalization.
+Additionally, robotic embodiment heterogeneity Wang et al.
+(2024) causes significant variation in VLA outputs, making end-to-end supervised learning from vision and language to diverse action outputs a potentially unrealistic path toward generalization.
+We present a novel pipeline addressing generalization challenges in robotic manipulation.
 Our approach leverages VLMs’ visual understanding capabilities, augmented with step-by-step reasoning to extract unified mid-level structural representations independent of robot embodiment.
@@
 These visual aids provide expressive yet compact spatial information that enables effective reasoning and decisions, overcoming both scarcity and heterogeneity limitations.
-We introduce FSD (From Seeing to Doing), a model that generates visual intermediate representations through spatial understanding and reasoning (Fig.
+We introduce FSD (From Seeing to Doing), a model that generates visual intermediate representations through spatial reasoning (Fig.
 1).
-The FSD model comprises three key components: (1) Spatial Relationship-Focused Visual Chain-of-Thought (SrCoT), which conducts multi-step reasoning anchored by object coordinates and spatial relationships, treating visual aid generation as a reasoning process; (2) A hierarchical data pipeline combining large-scale embodied datasets with common sense data, establishing a weak-to-strong capability enhancement training process; and (3) A self-consistency mechanism that aligns understanding and generation by binding spatial coordinates with specific visual signals.
-We employ supervised fine-tuning of VLMs using annotated spatial COT datasets for visual aids generation, then action executors follow spatial affordances and visual traces through simple planning methods to complete action execution.
-Our FSD model generalizes effectively to new instructions, and scenes through enhanced reasoning abilities.
-Our contributions include: (1) A novel paradigm bridging VLM reasoning with embodied decisions via visual aids; (2) The SrCoT method enabling multi-step reasoning for visual aid generation and guiding zero-shot manipulation; (3) Our weak-to-strong spatial reasoning and visual aids datasets, along with VABench, a manually annotated challenging benchmark for visual aids generation; (4) Superior performance across 8 benchmarks in spatial reasoning, free space reference, and visual aids generation, with zero-shot deployment achieving 54.1% success in SimplerEnv and 73% in 8 real-world tasks, outperforming mainstream VLA models by 30%.
+FSD comprises three key components: (1) Spatial Relationship-Focused Visual Chain-of-Thought (SrCoT), which conducts multi-step reasoning anchored by object coordinates and spatial relationships, treating visual aid generation as a reasoning process; (2) A hierarchical data construction pipeline combining large-scale embodied datasets with common sense data, establishing a weak-to-strong capability enhancement training process; and (3) A self-consistency mechanism that aligns understanding and generation by binding spatial coordinates with specific visual signals.
+To evaluate the accuracy and generalization ability of visual aids generated in complex scenes, we also manually annotated 300 images from both real-world and simulation tasks across various scenarios, forming a Visual Aids Generation Benchmark (VABench).
+Through carefully crafted training and evaluation methods, FSD can achieve precise generation of visual aids in different scenarios, then the robot follows spatial affordances and visual traces through simple planning methods to complete action execution.
+FSD generalizes effectively to new instructions, and scenes through enhanced reasoning abilities.
+Our contributions include: (1) A novel paradigm bridging VLM reasoning with embodied decisions via visual aids; (2) The SrCoT method enabling multi-step reasoning for visual aid generation and guiding zero-shot manipulation; (3) Our weak-to-strong spatial reasoning and visual aids datasets, along with VABench, a manually annotated challenging benchmark for visual aids generation; (4) Superior performance across 8 benchmarks in spatial reasoning, free space reference, and visual aids generation, with zero-shot deployment achieving 40.6% success in SimplerEnv and 72% in 8 real-world tasks, outperforming RoboPoint baseline by 30%.
 2 Related Work
-Spatial Understanding and Reasoning with VLMs Spatial understanding and reasoning [32, 48, 8, 28, 44] require VLMs to infer spatial information beyond 2D RGB images, a capability crucial for embodied AI applications such as navigation [48, 12, 22] and manipulation [66].
-Recent advances include SpatialVLM [5], which explicitly incorporates spatial primitives and coordinate systems to enhance geometric reasoning capabilities.
-Similarly, SpatialRGPT [7] and SpatialBot [4] improve spatial capabilities through more precise spatial relationship modeling.
+Spatial Understanding and Reasoning with VLMs Spatial understanding and reasoning Liu et al.
+(2023b), Song et al.
+(2024), Du et al.
+(2024), Liao et al.
+(2024), Ray et al.
+(2024) require VLMs to infer spatial information beyond 2D RGB images, a capability crucial for embodied AI applications such as navigation Song et al.
+(2024), Hong et al.
+(2023), Li et al.
+(2024b) and manipulation Yuan et al.
+(2024b).
+Recent advances include SpatialVLM Chen et al.
+(2024), which explicitly incorporates spatial primitives and coordinate systems to enhance geometric reasoning capabilities.
+Similarly, SpatialRGPT Cheng et al.
+(2024) and SpatialBot Cai et al.
+(2024) improve spatial capabilities through more precise spatial relationship modeling.
 FSD specifically targets embodied manipulation scenarios by enhancing spatial reasoning capabilities through novel SrCoT mechanisms and sophisticated self-consistency alignment techniques.
-Visual Chain-of-thought Reasoning The integration of Chain-of-thought (CoT) [56] and its variants [69, 63, 62] has significantly enhanced LLM reasoning abilities through structured step-by-step processes.
-For the multimodal reasoning challenge, researchers have developed various CoT approaches [37, 70, 61, 58] that establish appropriate reasoning anchors and extend reasoning pathways.
-Shikra [6] improves referential expression by incorporating specific visual regions into the reasoning process, while VoCoT [27] extends visually-grounded reasoning chains.
-VisualCoT [46] provides a comprehensive benchmark validating CoT effectiveness in multi-hop image reasoning tasks.
-EmbodiedCoT [67] pioneered CoT application in embodied AI by enhancing intermediate reasoning through fine-tuning OpenVLA [21].
+Visual Chain-of-thought Reasoning The integration of Chain-of-thought (CoT) Wei et al.
+(2022) and its variants Zhang et al.
+(2022), Yao et al.
+(2023b, 2024) has significantly enhanced LLM reasoning abilities through structured step-by-step processes.
+For the multimodal reasoning challenge, researchers have developed various CoT approaches Mitra et al.
+(2024), Zheng et al.
+(2023a), Yao et al.
+(2023a), Wu et al.
+(2025) that establish appropriate reasoning anchors and extend reasoning pathways.
+Shikra Chen et al.
+(2023) improves referential expression by incorporating specific visual regions into the reasoning process, while VoCoT Li et al.
+(2024f) extends visually-grounded reasoning chains.
+VisualCoT Shao et al.
+(2024) provides a comprehensive benchmark validating CoT effectiveness in multi-hop image reasoning tasks.
+EmbodiedCoT Zawalski et al.
+(2024) pioneered CoT application in embodied AI by enhancing intermediate reasoning through fine-tuning OpenVLA Kim et al.
+(2024).
 In contrast, FSD uniquely integrates visual-spatial reasoning using spatial relationship graphs as reasoning anchors.
 Visual Aids Empowered Robotic Manipulation Extracting embodiment-agnostic visual aids to enhance training efficiency has emerged as a promising paradigm in robotic manipulation.
-Numerous studies [1, 57, 59, 72, 65] have explored robotic policy learning based on visual traces, though most remain confined to specific tasks with cross-embodiment applicability.
-LLaRVA [39] advances this field by predicting visual traces to better align visual and action spaces for robot learning, compiling an impressive large visual trace dataset, yet struggles to generalize to novel downstream tasks without task-specific fine-tuning.
-Spatial affordance represents another effective visual aid, with several works [66, 38, 42, 48, 16, 60, 26] demonstrating its utility in robotic manipulation tasks.
+Numerous studies Bharadhwaj et al.
+(2024), Wen et al.
+(2023), Xu et al.
+(2024), Zheng et al.
+(2024), Yuan et al.
+(2024a) have explored robotic policy learning based on visual traces, though most remain confined to specific tasks with cross-embodiment applicability.
+LLaRVA Niu et al.
+(2024) advances this field by predicting visual traces to better align visual and action spaces for robot learning, compiling an impressive large visual trace dataset, yet struggles to generalize to novel downstream tasks without task-specific fine-tuning.
+Spatial affordance represents another effective visual aid, with several works Yuan et al.
+(2024b), Mo et al.
+(2021), Qin et al.
+(2020), Song et al.
+(2024), Ji et al.
+(2025), Yang et al.
+(2025), Li et al.
+(2025) demonstrating its utility in robotic manipulation tasks.
 Our key insight is that the scarcity of embodied data fundamentally limits purely data-driven visual aid prediction approaches, impeding zero-shot generalization to unseen scenarios.
@@
 These traces enable complex instruction execution, cross-embodiment transfer, and collision avoidance.
-We implement these representations in 2D rather than 3D due to limited high-quality 3D data availability [68].
+We implement these representations in 2D rather than 3D due to limited high-quality 3D data availability Zhang et al.
+(2024).
 Our object-centric rather than agent-centric visual aids approach effectively circumvents the limitations of heterogeneous embodied data while leveraging general visual datasets without robots, thus enabling robust generalization to novel scenarios and tasks.
@@
 3.2 Spatial Relationship-Focused Visual Chain-of-thought
-To enable VLMs to generate spatial visual aids, a direct approach is supervised fine-tuning [39, 66, 23, 67] using these aids as a new action space or employing generative models [47, 59].
+To enable VLMs to generate spatial visual aids, a direct approach is supervised fine-tuning Niu et al.
+(2024), Yuan et al.
+(2024b), Li et al.
+(2024c), Zawalski et al.
+(2024) using these aids as a new action space or employing generative models Shridhar et al.
+(2024), Xu et al.
+(2024).
 However, the heterogeneity and scarcity of embodied datasets limit this method.
@@
 This step-by-step SrCoT approach, though powerful, fundamentally depends on precise spatial understanding capabilities.
+To improve the stability and reliability of reasoning paths and reduce model hallucinations, SrCoT requires the model to generate coordinates in a specified format and bind them to objects while performing object-centric reasoning Wang et al.
+(2025), Li et al.
+(2024f).
+We use <ref> to mark objects, and <point> and <box> to mark points and boxes, respectively, ensuring that each object is strictly bound to its corresponding coordinates.
+This explicit visual-spatial coordinate alignment enhances the FSD ’s understanding of the spatial positions and relationships of objects.
+All coordinates are treated as text and are discretized as integers normalized between 0 and 999.
 3.3 Weak-to-Strong Capability Dataset Construction Pipeline
-The SrCoT mechanism demands enhanced capabilities from VLMs, including precise reference grounding, spatial understanding, and complex instruction-following capabilities (directly predicting future point trajectories) where mainstream VLMs [34, 30, 44, 8] show limitations.
+The SrCoT mechanism demands enhanced capabilities from VLMs, including precise reference grounding, spatial understanding, and complex instruction-following capabilities (directly predicting future point trajectories) where mainstream VLMs Liu et al.
+(2024b), Lin et al.
+(2024b), Ray et al.
+(2024), Du et al.
+(2024) show limitations.
 Consequently, we designed a weak-to-strong data construction pipeline to progressively develop these abilities.
 FSD encompasses five hierarchical capability levels: 1 Region Grounding enables robots to focus on key objects in scenes.
-Although grounding capabilities have been broadly integrated into current VLMs [6, 64], understanding various small objects and complex scenes for embodied tasks is still limited; 2 Spatial Relationship understanding establishes prerequisite knowledge for spatial reasoning, forming the anchor points for SrCoT; 3 Spatial Reasoning builds upon these foundations to perform multi-hop analysis of object positions and relationships; and finally, 4 Spatial Affordance Generation and 5 Visual Trace Generation create actionable spatial guidance.
+Although grounding capabilities have been broadly integrated into current VLMs Chen et al.
+(2023), You et al.
+(2023), understanding various small objects and complex scenes for embodied tasks is still limited; 2 Spatial Relationship understanding establishes prerequisite knowledge for spatial reasoning, forming the anchor points for SrCoT; 3 Spatial Reasoning builds upon these foundations to perform multi-hop analysis of object positions and relationships; and finally, 4 Spatial Affordance Generation and 5 Visual Trace Generation create actionable spatial guidance.
 Notably, SrCoT functions as a general visual-spatial reasoning mechanism applicable beyond visual traces to diverse spatial reasoning tasks.
@@
 Figure 4: FSD screens data from large-scale embodied datasets, generates ground truth spatial relationship graph. We finally collected 300K data for 10+ embodiments with 5-level capabilities.
-Automatic Dataset Construction: We leveraged extensive robot data from BridgeDataV2 [53], RT-X [40], and Droid [20] to construct FSD’s training dataset.
-Inspired by SpatialVLM [5] and SpatialRGPT [7], our automated data construction pipeline created 300K Supervised Fine-Tuning (SFT) data across numerous formats.
-After filtering demonstrations with unclear instructions, we used GPT4o [15] to nominate task-relevant objects while excluding out-of-range or overly complex items.
-We then built around objects using GroundedSam [45] for bounding boxes and segmentation masks (Level 1 dataset).
-For spatial relationship labels, we reconstructed 3D semantic scene graphs using Metric3Dv2 [13] for depth estimation, along with WildCamera [73] and PerspectiveFields [17] for camera parameters.
+Automatic Dataset Construction: We leveraged extensive robot data from BridgeDataV2 Walke et al.
+(2023), RT-X O’Neill et al.
+(2023), and Droid Khazatsky et al.
+(2024) to construct FSD’s training dataset.
+Inspired by SpatialVLM Chen et al.
+(2024) and SpatialRGPT Cheng et al.
+(2024), our automated data construction pipeline created 300K Supervised Fine-Tuning (SFT) data across numerous formats.
+After filtering demonstrations with unclear instructions, we used GPT4o Hurst et al.
+(2024) to nominate task-relevant objects while excluding out-of-range or overly complex items.
+We then built around objects using GroundedSam Ren et al.
+(2024) for bounding boxes and segmentation masks (Level 1 dataset).
+For spatial relationship labels, we reconstructed 3D semantic scene graphs using Metric3Dv2 Hu et al.
+(2024) for depth estimation, along with WildCamera Zhu et al.
+(2024) and PerspectiveFields Jin et al.
+(2023) for camera parameters.
 This enabled 2D-3D mapping and spatial relationship graph construction.
@@
 To create spatial affordance labels (Level 4 Dataset), we extract the manipulated object’s final position from the terminal frame, combine it with reference object positioning, calculate the precise affordance region, and re-render this information onto the initial frame.
-For visual trace generation (Level 5 Dataset), we employ a two-stage approach: first applying self-supervised keypoint extraction [14] to identify grasp points on manipulated objects, then utilizing Cotracker [19] to capture temporal dynamics from human demonstrations, subsequently projecting these trajectories onto the initial frame.
+For visual trace generation (Level 5 Dataset), we employ a two-stage approach: first applying self-supervised keypoint extraction Huang et al.
+(2024) to identify grasp points on manipulated objects, then utilizing Cotracker Karaev et al.
+(2024) to capture temporal dynamics from human demonstrations, subsequently projecting these trajectories onto the initial frame.
 Throughout this process, we employed strict rule-based filters and continually validated our approach against manually annotated test sets, iteratively refining our filtering criteria based on empirical feedback to ensure the resulting dataset met our quality requirements.
@@
 3.4 Self-Consistent Alignment for Spatial Understanding and Generation
-High-quality SFT datasets enable VLMs to generate visual aids [66], yet these models struggle to understand the physical meaning of such annotations since coordinate spaces never appeared in pretraining data.
+High-quality SFT datasets enable VLMs to generate visual aids Yuan et al.
+(2024b), yet these models struggle to understand the physical meaning of such annotations since coordinate spaces never appeared in pretraining data.
 The alignment between image coordinates and actual spatial positions presents a significant challenge.
@@
 4 Training and Action Execution of FSD
-Training: We follow the instruction tuning pipeline proposed by LLaVA-1.5 [33, 34].
+Training: We follow the instruction tuning pipeline proposed by LLaVA-1.5 Liu et al.
+(2023c, 2024b).
 As shown in Fig.
-1, FSD’s architecture comprises an image encoder (CLIP-ViT-L-336px [11]), a linear projector, a language tokenizer, and a LLM (Vicuna-13B [71]).
+1, FSD’s architecture comprises an image encoder (CLIP-ViT-L-336px Gao et al.
+(2024)), a linear projector, a language tokenizer, and a LLM (Vicuna-13B Zheng et al.
+(2023b)).
 The image encoder processes images into tokens.
@@
 Only the projector and LLM weights are updated during fine-tuning while the vision encoder and tokenizer remain frozen.
-We built upon ASMv2 [55] as our foundation, which already incorporates basic relation conversation and reference grounding capabilities.
+We built upon ASMv2 Wang et al.
+(2025) as our foundation, which already incorporates basic relation conversation and reference grounding capabilities.
 The training process of FSD is divided into two stages: General Spatial Reasoning Enhancement: Using data from levels 1-3, we focus on improving the model’s embodied spatial reasoning capabilities.
 Following Yuan et al.
-[66] and Brohan et al.
-[3], we discovered that an appropriate data mixture is crucial for downstream performance.
+(2024b) and Brohan et al.
+(2023), we discovered that an appropriate data mixture is crucial for downstream performance.
 Joint training with mixed robotic and internet data ensures the model retains knowledge acquired during pre-training.
@@
 Following the pinhole camera model, we employ depth-based back-projection to map these to 3D space, yielding $\boldsymbol{\tau}^{3d}=\{\mathbf{x}_{t}^{3d}\mid t=1,2,...,T\}$.
-Next, based on the spatial position of the first point $\mathbf{x}_{1}$, we query GraspNet’s [9] grasp candidates $G$ to match the nearest grasp pose $G^{*}$.
+Next, based on the spatial position of the first point $\mathbf{x}_{1}$, we query GraspNet’s Fang et al.
+(2020) grasp candidates $G$ to match the nearest grasp pose $G^{*}$.
 For relatively fixed scenes, we may also use predetermined grasp poses.
 Subsequently, we optimize the path trajectory using gradient descent-based interpolation, generating complete motion trajectories in SE(3) space, enabling the robotic arm to follow the 3D visual trajectory.
-When using only spatial affordance, we utilize CuRobo [49] as the motion planner to determine execution trajectories $T$ based on the target position $\mathcal{P}$.
+When using only spatial affordance, we utilize CuRobo Sundaralingam et al.
+(2023) as the motion planner to determine execution trajectories $T$ based on the target position $\mathcal{P}$.
 More details are provided in Appendix C.
-Unlike methods such as LLARVA [39] and EmbodiedCOT [67] which also utilize visual auxiliary aids, FSD transforms prediction tasks into reasoning tasks, better leveraging visual-spatial common knowledge without requiring scenario-specific fine-tuning.
+Unlike methods such as LLARVA Niu et al.
+(2024) and EmbodiedCOT Zawalski et al.
+(2024) which also utilize visual auxiliary aids, FSD transforms prediction tasks into reasoning tasks, better leveraging visual-spatial common knowledge without requiring scenario-specific fine-tuning.
 5 Visual Aids Generation Benchmark
-Few datasets exist for evaluating visual aid generation, Where2Place [66] provides 100 real-world images from homes and offices but is limited to direct and simple language instructions.
+Few datasets exist for evaluating visual aid generation, Where2Place Yuan et al.
+(2024b) provides 100 real-world images from homes and offices but is limited to direct and simple language instructions.
 And no benchmarks for trajectory prediction.
 To address this gap, we propose the Visual Aids Generation Benchmark (VABench).
-we manually annotated 300 problems from real-world and simulation datasets (OXE [40], BridgeData [53], Droid [20], Libero [31]), requiring models to infer visual aids given only natural language instructions similar to everyday human commands.
+we manually annotated 300 problems from real-world and simulation datasets (OXE O’Neill et al.
+(2023), BridgeData Walke et al.
+(2023), Droid Khazatsky et al.
+(2024) and Libero Liu et al.
+(2023a)), requiring models to infer visual aids given only natural language instructions similar to everyday human commands.
 For spatial affordance (VABench-Point), we measure the proportion of points falling within target regions.
@@
 We interpolate when trajectory lengths differ and normalize coordinates to a 1000×1000 space for consistent evaluation.
-Since multiple valid solutions exist for each instruction, we simulated human evaluation standards by establishing detailed assessment criteria and added metrics with MLLM-based qualitative scoring (0-10) of visualized trajectories, named GPT Score.
+Since multiple valid solutions exist for each instruction, we simulated human evaluation standards by establishing detailed assessment criteria and added metrics with MLLM-based qualitative scoring (1-10) of visualized trajectories, named GPT Score.
 We provide detailed dataset information and evaluation procedures in Appendix D.
@@
 For Seeing, we tested its general spatial reasoning and visual aids generation capabilities.
-For Doing, we conducted zero-shot manipulation experiments in both SimplerEnv [25] simulation and real-world xArm robotic platforms to assess its practical generalization performance.
+For Doing, we conducted zero-shot manipulation experiments in both SimplerEnv Li et al.
+(2024e) simulation and real-world xArm robotic platforms to assess its practical generalization performance.
 6.1 Evaluation of Spatial Understanding and Reasoning Capabilities
 Benchmarks and Baselines.
-General Spatial Reasoning Capabilities: We evaluated general spatial reasoning capabilities using five popular benchmarks: CVBench [51], BLINK [10], CRPE [55], SAT [44], and EmbSpatial-Bench [8].
+General Spatial Reasoning Capabilities: We evaluated general spatial reasoning capabilities using five popular benchmarks: CVBench Tong et al.
+(2024), BLINK Fu et al.
+(2024), CRPE Wang et al.
+(2025), SAT Ray et al.
+(2024), and EmbSpatial-Bench Du et al.
+(2024).
 These benchmarks encompass 15 subtasks measuring various spatial competencies including counting, spatial relationships, distance estimation and so on.
-We included two leading closed-source models: GPT-4o [15] and GPT-4V as performance reference.
-Subsequently, we conducted comparative analyses against other open-source spatial enhanced MLLMs, including LLaVA-1.5 [33], SAT-Dynamic [44], RoboPoint [66], and ASMv2 [55], all with 13B parameters.
+We included two leading closed-source models: GPT-4o Hurst et al.
+(2024) and GPT-4V as performance reference.
+Subsequently, we conducted comparative analyses against other open-source spatial enhanced MLLMs, including LLaVA-1.5 Liu et al.
+(2023c), SAT-Dynamic Ray et al.
+(2024), RoboPoint Yuan et al.
+(2024b), and ASMv2 Wang et al.
+(2025), all with 13B parameters.
 Object and Free Region Reference Capabilities: Following Yuan et al.
-[66], we assessed embodied spatial capabilities using the RoboRefIt [36] and Where2Place [66] benchmarks.
-We compared mainstream closed-source models and MLLM for enhancing spatial abilities (SpatialBot [4], SpaceLLaVA [5], and RoboBrain [16]).
+(2024b), we assessed embodied spatial capabilities using the RoboRefIt Lu et al.
+(2023) and Where2Place Yuan et al.
+(2024b) benchmarks.
+We compared mainstream closed-source models and MLLM for enhancing spatial abilities (SpatialBot Cai et al.
+(2024), SpaceLLaVA Chen et al.
+(2024), and RoboBrain Ji et al.
+(2025)).
 We used the proportion of predicted points within specified objects/regions as the accuracy metric.
@@
 Spatial Affordance and Visual Trace Capabilities: We utilized our VABench to evaluate the capabilities.
-We found very few models with this capability for Visual Trace prediction, so we trained an end-to-end prediction baseline model using a pre-trained DINOv2 [41] encoder coupled with a multi-layer transformer [52] architecture to predict visual trajectories, trained on the same visual trajectory data, named DINOv2 Predictor.
+We found very few models with this capability for Visual Trace prediction, so we trained an end-to-end prediction baseline model using a pre-trained DINOv2 Oquab et al.
+(2023) encoder coupled with a multi-layer transformer Vaswani et al.
+(2017) architecture to predict visual trajectories, trained on the same visual trajectory data, named DINOv2 Predictor.
 We conducted this comparison to demonstrate the advantages of our reasoning-based FSD approach.
@@
 See Appendix F for more visualizations.
+Figure 6: Visualization of FSD executing tasks. The first six rows are real-world experiments, and the last four rows are from SIMPLEREnv.
+Figure 7: An example and visualization of FSD for generating visual trace.
+Figure 8: An example and visualization of FSD for generating affordance box and points.
 Table 1: Performance comparison on 5 spatial reasoning benchmarks. Bold and underlined values show best and second-best performance among open-source models.
-CVBench CRPE SAT BLINK EmbSpatial Rank
+CVBench CRPE SAT BLINK EmbSp Rank
 Count 2DRel 3DDep 3DDis Avg. Exist. Subj. Pred. Obj. Avg. Val Real Count MV RelDepth SpRel Avg. Test
-Closed-source models
+Closed-source
 GPT-4V 62.4 71.1 79.8 68.3 70.4 90.6 76.7 65.1 68.5 75.2 44.8 50.7 60.8 55.6 59.7 72.7 62.2 36.1 -
 GPT-4o 65.9 85.5 87.8 78.2 79.4 93.3 81.9 71.8 73.6 80.2 49.4 57.5 49.2 60.2 74.2 69.2 63.2 49.1 -
-Open-source models
+Open-source
 LLaVA-1.5-13B 58.2 46.6 53.0 47.8 51.4 88.7 57.4 54.2 55.2 63.9 51.4 41.6 45.0 41.4 53.2 69.9 52.4 35.1 4.8
@@
 […]
-DINOv2 Predictor 128.32 117.49 2.17
-FSD 78.26 63.44 5.49
-   w/o SrCoT 99.53 80.06 4.00
-   w/o Alignment 80.48 66.80 5.45
-Table 4: SimplerEnv Evaluation on WidowX Robot The results of baselines are derived from [43]. ZS: zero-shot, FT: fine-tuning using BridgeData. Each task is tested 24 episodes.
-Model Spoon→Towel Carrot→Plate Green→Yellow Eggplant→Basket Avg.
+GPT4o 136.13 113.53 4.37
+DINOv2 Predictor 128.32 117.49 4.01
+FSD 78.26 63.44 6.21
+   w/o SrCoT 99.53 80.06 5.07
+   w/o Alignment 80.48 66.80 5.92
+Table 4: SimplerEnv Evaluation on WidowX Robot. The results of baselines are derived from Qu et al. (2025). ZS: zero-shot, FT: fine-tuning using BridgeData. Each task is tested 24 episodes.
+[…]
 Grasp Succ. Grasp Succ. Grasp Succ. Grasp Succ.
-RT-1-X [40] 16.7% 0% 20.8% 4.2% 8.3% 0% 0.0% 0% 1.1%
-Octo-S [50] 77.8% 47.2% 27.8% 9.7% 40.3% 4.2% 87.5% 56.9% 30.0%
-OpenVLA [21] 4.1% 0% 33.3% 0% 12.5% 0% 8.3% 4.1% 1.0%
-RoboVLM (ZS) [24] 37.5% 20.8% 33.3% 25.0% 8.3% 8.3% 0.0% 0% 13.5%
-RoboVLM (FT) [24] 54.2% 29.2% 25.0% 25.0% 45.8% 12.5% 58.3% 58.3% 31.3%
-SpatialVLA (ZS) [43] 25.0% 20.8% 41.7% 20.8% 58.3% 25.0% 79.2% 70.8% 34.4%
-SpatialVLA (FT) [43] 20.8% 16.7% 29.2% 25.0% 62.5% 29.2% 100% 100% 42.7%
-FSD 58.3% 41.6% 58.3% 50.0% 91.6% 33.3% 37.5% 37.5% 54.1%
+RT-1-X O’Neill et al. (2023) 16.7% 0% 20.8% 4.2% 8.3% 0% 0.0% 0% 1.1%
+Octo-S Team et al. (2024) 77.8% 47.2% 27.8% 9.7% 40.3% 4.2% 87.5% 56.9% 30.0%
+OpenVLA Kim et al. (2024) 4.1% 0% 33.3% 0% 12.5% 0% 8.3% 4.1% 1.0%
+RoboVLM (ZS) Li et al. (2024d) 37.5% 20.8% 33.3% 25.0% 8.3% 8.3% 0.0% 0% 13.5%
+RoboVLM (FT) Li et al. (2024d) 54.2% 29.2% 25.0% 25.0% 45.8% 12.5% 58.3% 58.3% 31.3%
+SpatialVLA (ZS) Qu et al. (2025) 25.0% 20.8% 41.7% 20.8% 58.3% 25.0% 79.2% 70.8% 34.4%
+SpatialVLA (FT) Qu et al. (2025) 20.8% 16.7% 29.2% 25.0% 62.5% 29.2% 100% 100% 42.7%
+FSD 58.3% 41.7% 58.3% 50.0% 91.7% 33.3% 37.5% 37.5% 40.6%
 FSD demonstrates breakthrough performance in visual aid generation.
@@
 Ablation studies confirm the critical contributions of both SrCoT and self-consistency alignment, validating our reasoning-based approach where step-by-step spatial analysis enables more accurate predictions than purely data-driven methods.
-We provide detailed comparisons with RoboBrain [16] in Appendix I.
+We provide detailed comparisons with RoboBrain Ji et al.
+(2025) in Appendix I.
 Visualization of visual aids generated by FSD We present examples of generation in Fig.
@@
 We also presented the reasoning process of FSD in Fig.
-6, where the CoT provides clear guidance for the results.
+9, where the CoT provides clear guidance for the results.
+Subsequently, we selected several specific examples to showcase the complete outputs of FSD in Fig.
+8 and Fig.
+7.
 More visualizations are in Appendix F.
-6.2 Evaluate the Decision Capability
-SimplerEnv Simulation SimplerEnv [25] is a simulation environment specifically designed for evaluating real-world robotic manipulation.
+6.2 Evaluation of the Decision Capability
+SimplerEnv Simulation SimplerEnv Li et al.
+(2024e) is a simulation environment specifically designed for evaluating real-world robotic manipulation.
 It provides a standardized platform, emphasizing reproducibility with real-world scenarios.
 We utilized FSD to generate visual traces and perform zero-shot deployment on the WidowX robotic arm.
-We compared its performance against state-of-the-art general manipulation VLA, including Octo [50], OpenVLA [21], RoboVLM [24] and SpatialVLA [43].
+We compared its performance against state-of-the-art general manipulation VLA, including Octo Team et al.
+(2024), OpenVLA Kim et al.
+(2024), RoboVLM Li et al.
+(2024d) and SpatialVLA Qu et al.
+(2025).
 The results are presented in Table 4.
-FSD surpasses the baseline in most tasks without FT with a 54.1% success rate, demonstrating the strong generalization ability of FSD.
-Notably, FSD achieves the best average performance, even outperforming VLA models that even outperforming specially fine-tuned VLA models.
+FSD surpasses the baseline in most tasks without FT with a 40.6% success rate, demonstrating the strong generalization ability of FSD.
+Notably, FSD achieves the best zero-shot average performance and is competitive with specially fine-tuned VLA models.
 This superior generalization capability demonstrates that visual aids representations provide better zero-shot transfer compared to direct action prediction approaches, bridging the gap between perception and control more effectively.
 Real World Robot Evaluation As shown in Fig.
-7, we conducted zero-shot tests with FSD on an xArm 6 robot across 8 tabletop manipulation tasks.
+10, we conducted zero-shot tests with FSD on an xArm 6 robot across 8 tabletop manipulation tasks.
 The setup included an Intel RealSense L515 LiDAR camera.
@@
 Full results are presented in Fig.
-7 and we refer to Appendix G for detailed setup and more visualizations.
+10 and visualization of control tasks executed using FSD in Fig.
+6.
+We refer to Appendix G for detailed setup.
 7 Conclusion
@@
 Limitations include reliance on 2D rather than 3D trajectory generation and constraints from available training data quality.
-Future works are detailed in Appendix J.
+More limitations and future works are in Appendix J.
+8 Acknowledgment
+We would like to thank Zhongwen Xu, Liang Wang, Shuyang Gu, and Chen Li for their participation in the discussions of this paper and for providing valuable insights.
+In addition, we would especially like to thank Yiyang Huang for the constructive suggestions on improving the figures in the manuscript.
+Appendix
+Appendix A Weak-to-Strong Dataset Construction
+Region Grounding Dataset Generation Regional grounding enables robots to focus on key objects within a scene.
+In embodied AI, robots need to concentrate on task-relevant local scenes in the image according to task instructions and be able to accurately locate objects.
+In Level 1 of our dataset, we integrate grounding with captioning, aiming for the agent to provide positional information when describing images.
+Unlike traditional image captioning, objects in embodied scenes are often cluttered, and we only seek to extract captions for task-relevant regions, avoiding redundant information.
+We prompt GPT-4o to exclude task-irrelevant regions based on task instructions and generate both image captions and the corresponding object names.
+Subsequently, we use Grounding DINO Liu et al.
+(2023d) to capture the locations of various objects in the image and embed them into the caption sentences, forming captions with object location information.
+Below is an example from the dataset:
+Level 1: Region Grounding Human: Give a short and clear explanation of the subsequent image.
+GPT: A <ref>robotic arm</ref><box>[[338, 126, 994, 861]]</box> is positioned above a <ref>blue bottle</ref><box>[[257, 201, 381, 413]]</box>, with a <ref>grey toy</ref><box>[[391, 413, 538, 518]]</box> to the left and a <ref>green oval object</ref><box>[[592, 481, 702, 601]]</box> to the right on a wooden table.
+Spatial Relationship Dataset Generation To accurately infer object spatial relationships from RGB images, a multi-stage pipeline is employed: encompassing object detection, instance segmentation, 2D-to-3D mapping, and subsequent relationship calculation.
+Initially, for pre-identified objects of interest within the scene, GroundedSAM Ren et al.
+(2024) is utilized to perform instance segmentation, yielding precise object masks.
+Subsequently, the 2D RGB information is elevated to a 3D spatial representation.
+This transformation begins with leveraging PerspectiveFields Jin et al.
+(2023) to estimate the Z-axis orientation, serving as a coarse approximation for camera extrinsics.
+Concurrently, the WildCamera Zhu et al.
+(2024) model is employed to estimate intrinsic camera parameters, including focal length and resolution.
+Metric3Dv2 Hu et al.
+(2024) is then used for robust depth estimation.
+By integrating the RGB image, estimated depth image, and computed camera intrinsics and extrinsics, the 2D RGB image is transformed into a 3D point cloud.
+Given the prior object segmentation, the specific 3D spatial position and size of each individual object can be precisely derived from the generated point cloud.
+These comprehensive 3D data then enable the calculation of relative spatial relationships between objects, which are subsequently exported as a spatial relationship graph.
+We present the 2D RGB-D images before transformation and the visualized point clouds after transformation in Fig.
+11.
+Figure 11: RGB-D images before transformation and the visualized point clouds after transformation.
+It is important to note that for inferring spatial relationships, only relative depth sorting data is generated, which means stringent accuracy requirements for absolute depth are relaxed.
+To further enhance data quality, particularly for robust relationship inference, objects exhibiting a relative depth gap of at least 20% are preferentially selected for subsequent generation.
+An example from the dataset:
+Level 2: Spatial Relationship Human: Tell me about the items in your view and their relationships to each other. Answer the question with scene graph.
+GPT: <ref>front left burner</ref><box>[[597, 422, 817, 596]]</box> is positioned <pred>to the right of</pred><box>[[597, 422, 817, 596]]</box><box>[[171, 275, 446, 409]]</box> <ref>metal pot</ref><box>[[171, 275, 446, 409]]</box>.
+<ref>metal pot</ref><box>[[171, 275, 446, 409]]</box> appears <pred>to the left of</pred><box>[[171, 275, 446, 409]]</box><box>[[431, 474, 715, 608]]</box> <ref>green spatula</ref><box>[[431, 474, 715, 608]]</box>.
+<ref>metal pot</ref><box>[[171, 275, 446, 409]]</box> can be seen <pred>to the left of</pred><box>[[171, 275, 446, 409]]</box><box>[[519, 408, 600, 490]]</box> <ref>orange cheese slice</ref><box>[[519, 408, 600, 490]]</box>.
+Spatial Reasoning QA Generation After generating the spatial relation diagrams, we can easily create various template questions for spatial reasoning Cheng et al.
+(2024), Du et al.
+(2024), such as: How are [A] and [B] positioned in relation to each other in the image?
+or From your perspective, which object in the image is at the shortest distance?
+In addition to template-based QA, we also combine task instructions, existing spatial information, and images to query GPT-4o, thereby generating more diverse multi-turn dialogues to enhance the model’s generalization ability in spatial reasoning.
+Spatial Affordance and Visual Trace Dataset Generation Next, we provide a detailed description of how to extract the required controllable points/boxes and visual trajectories from the embodied dataset like BridgeData Walke et al.
+(2023), RT-X O’Neill et al.
+(2023) and Droid Khazatsky et al.
+(2024).
+Our methodology for extracting visual aids, specifically Affordance Boxes, Affordance Points, and Visual Traces, involves a multi-stage process leveraging state-of-the-art vision models.
+We also incorporate a rigorous data validation procedure to ensure high-quality output.
+First, we acquire the initial and final frames of the video sequence.
+To determine the Affordance Box, we utilize GroundingDINO Liu et al.
+(2023d) and GroundedSam Ren et al.
+(2024) to detect the mask of the manipulated_object in the final frame.
+The bounding box of this detected object in the final frame defines the Affordance Box, representing the ultimate spatial location of the manipulated object.
+Subsequently, we extract Affordance Points.
+This involves performing an erosion operation on the mask of the manipulated_object obtained from the final frame.
+Erosion reduces the mask’s area, facilitating the sampling of points that are more central or internal to the object, thereby mitigating the risk of sampling edge points.
+From this eroded mask, we uniformly sample 8 points, which constitute the Affordance Points.
+For Visual Trace extraction, we begin by detecting the mask of the manipulated_object in the first frame.
+From this mask, we sample 3 points, which serve as the initial query points for the CoTracker Karaev et al.
+(2024) model.
+Selecting multiple points enhances the robustness of trajectory tracking.
+The processed video sequence and these sampled points are then fed into the CoTracker model, which outputs the predicted trajectory for each query point across every frame of the video sequence.
+We calculate the total distance of each trajectory and select the longest trajectory as the representative trajectory for the manipulated_object.
+This chosen trajectory is then subjected to cubic spline interpolation for smoothing.
+Finally, 8 equidistant points are uniformly sampled from the smoothed trajectory, forming the Visual Trace.
+A critical aspect of this process is addressing potential prediction errors from the pre-trained visual models, such as incorrect object identification or incomplete tracking of object motion.
+To mitigate these issues, we implement stringent rule-based filtering using hyperparameters like size thresholds and trajectory length thresholds.
+Before annotating each dataset, we iteratively adjust these rules and conduct manual inspections on 100 examples.
+Only when the accuracy of these filtered results surpasses 95% do we deem the filtering rules robust enough to proceed with the full data generation pipeline.
+This meticulous validation process ensures the high quality of the generated data.
+Following the generation of these visual aids, we also pre-generate the thinking processes for Chain-of-Thought (CoT) reasoning.
+We input templates, questions, and answers into GPT-4o, querying it to complete the thinking process.
+The complete query prompt is constructed accordingly:
+Prompt and In-context Example for Completing Thinking Process You are an AI visual assistant that can analyze a single image.
+You receive one image and corresponding caption, task instruction, manipulated object in the task and target place to finish the task, and bounding box position and relative position of these objects.
+Also, you receive the answer of points and bbox.
+Now, you need to generate a <Description>...</Description>, <Reasoning>...</Reasoning>, <Answer>...</Answer> format.
+<Description> First, using the provided caption, task instruction, describe the scene.
+If there are errors in the caption, please ignore them and do not point them out in your description.
+Instead of directly mentioning the bounding box coordinates, utilize this data to explain the scene using natural language with its bounding box in the format like "<ref>object</ref><box>[[x1, y1, x2, y2]]</box>".
+When mentioning the predicate between two objects, you should mention it in the format like "<pred>predicate</pred><box>[[x1, y1, x2, y2]]</box><box>[[x3, y3, x4, y4]]</box>", where "<box>[[x1, y1, x2, y2]]</box>" denotes the bounding box coordinates of the subject and "<box>[[x3, y3, x4, y4]]</box>" denotes the bounding box coordinates of the object.
+</Description> <Reasoning> According to the task instruction and the answer of points and bbox, provide a chain-of-thought, logical explanation of the problem.
+</Reasoning> <Answer> State the final bounding box answer and point answer in a clear and direct format here.
+Bounding box answer is in the format like "<box>[[x1, y1, x2, y2]]</box>".
+Point answer is in the format like "<point>[[x1, y1], [x2, y2], [x3, y3], ...]</point>".
+</Answer> Here is an example:
+Caption: 7up can is on the table beside the bowl, drawer is open.
+Task instruction: place 7up can into top drawer
+Manipulated object: 7up can<box>[[686, 174, 788, 357]]</box>
+Target place: top drawer <box>[[135, 411, 852, 920]]</box> Current Relation: (7up can, [686, 174, 788, 357], top drawer, [135, 411, 852, 920], above)
+Bounding Box Answer: [569, 553, 651, 679]
+Point Answer: [[603, 609], [593, 637], [596, 643], [615, 596], [612, 593], [590, 637], [603, 606], [615, 631]]
+<Description> The scene describes a <ref>7up can</ref><box>[[686, 174, 788, 357]]</box> that is positioned on a table.
+It is situated <pred>above</pred><box>[[686, 174, 788, 357]]</box><box>[[135, 411, 852, 920]]</box> the <ref>top drawer</ref><box>[[135, 411, 852, 920]]</box> which is open.
+</Description> <Reasoning> The 7up can is above the top drawer now and should be placed into the top drawer.
+So the target position of the 7up can might be similar to the position of the drawer’s interior.
+</Reasoning> <Answer> To complete the task, the target position is <box>[[569, 553, 651, 679]]</box>.
+The new position can also be roughly defined by the following several free points <point>[[603, 609], [593, 637], [596, 643], [615, 596], [612, 593], [590, 637], [603, 606], [615, 631]]</point>, which can outline the position of the 7up can after it is placed in the top drawer.
+</Answer>
+Next, we present two examples from the dataset: one involves generating Affordance Boxes/Points (Level 4), and the other involves generating visual trace (Level 5).
+We consider Affordance Points as samples within the box; thus, generating the box first and then the points is a refinement process from coarse to fine granularity.
+Therefore, both types are generated for each data, and one can choose which to use as needed.
+In the task of generating visual trace, since we performed equidistant interpolation in advance, the visual trace in the dataset is fixed at 8 points.
+Consequently, models trained on such a dataset will also generate 8 points as the visual trace.
+Level 4: Spatial Affordance Box/Points Generation Human: "You are currently a robot performing robotic manipulation tasks.
+Your task instruction: Moved the can to the left side of the green cloth.
+Observe the image, use 2D points and bounding box to mark the target location where the manipulated object will be moved.
+In your answer, use <box>[[x1, y1, x2, y2]]</box> to present the bounding box of the target region, and use <point>[[x1, y1], [x2, y2], ...]</point> to mark the points of the free space." GPT: "<Description>
+The image shows a <ref>white soup can</ref><box>[[544, 177, 700, 368]]</box> positioned to the right of a <ref>green towel</ref><box>[[244, 389, 664, 676]]</box> on a table.
+A small bowl containing a yellow object and a brush is next to the towel.
+</Description>
+<Reasoning>
+The task is to move the white soup can to the left side of the green towel. Currently, the can is to the right of the towel, so it needs to be repositioned to the left of the towel’s current location.
+</Reasoning>
+<Answer>
+The target position for the white soup can is <box>[[250, 181, 400, 392]]</box>.
+The new position can also be roughly defined by the following points: <point>[[346, 248], [302, 365], [377, 251], [330, 295], [357, 291], [354, 362], [329, 355], [312, 352]]</point>.
+</Answer>"
+Level 5: Visual Trace Generation Human: "You are currently a robot performing robotic manipulation tasks.
+Your task instruction: put the green spatula in the silver pot.
+Observe the image, use 2D points to mark the manipulated object-centric waypoints to guide the robot to manipulate the object.
+Typically, the waypoints consist of an ordered sequence of eight 2D points.
+The format is <point>[[x1, y1], [x2, y2], ...]</point>." GPT: "<Description>
+The image shows a <ref>green spatula</ref><box>[[762, 536, 856, 711]]</box> positioned to the right of an orange towel on a table.
+A <ref>silver pot</ref><box>[[0, 723, 242, 872]]</box> is located above a robotic arm.
+Other objects, such as a gray can and a mushroom, are also present in the scene.
+</Description>
+<Reasoning>
+To move the green spatula into the silver pot, start by identifying the initial position of the spatula at <point>[[802, 613]]</point>.
+First, lift the spatula upwards to clear the table, moving to <point>[[774, 501]]</point>.
+Then, move it horizontally towards the left, passing through <point>[[685, 394]]</point>.
+Finally, lower the spatula into the silver pot, ending at <point>[[657, 401]]</point> within the target bounding box <box>[[225, 529, 407, 640]]</box>.
+</Reasoning>
+<Answer>
+The visual trace for placing the green spatula into the silver pot is <point>[[802, 613], [780, 582], [774, 501], [744, 465], [685, 394], [657, 349], [668, 354], [657, 401]]</point>.
+</Answer>"
+Appendix B Training Details and Datasets
+The training of FSD is a two-stage process, building upon the LLaVA Liu et al.
+(2024a) architecture through continued fine-tuning of ASMv2 Wang et al.
+(2025).
+In order to train a VLM with powerful spatiotemporal reasoning abilities that can also generate visual aids, we leveraged a comprehensive dataset of approximately 1.4M samples from various sources.
+Note: The position coordinate format used by FSD normalizes the image coordinates to a range of 0-999 after padding the image to a square shape.
+For all datasets mentioned below, the coordinates have been pre-processed in this manner.
+Stage 1: Spatial Reasoning Enhancement The primary goal of the first stage is to enhance the model’s spatial reasoning capabilities.
+This dataset is primarily composed of two parts: common-sense image QA and conversations data and spatial reasoning data.
+The inclusion of GeneralQA is crucial for FSD to maintain broad instruction-following abilities after fine-tuning.
+First, we selected some datasets and randomly sampled around 838k data from LLaVA-665k and ASMv2-4M instruction-following data.
+Next, we incorporated approximately 295k samples of spatial reasoning data from the LLaVA-OneVision Li et al.
+(2024a), RoboPoint Yuan et al.
+(2024b), SpatialBot Cai et al.
+(2024) and SAT Ray et al.
+(2024) training datasets.
+Finally, the first three levels of FSD’s collected embodied spatial reasoning datasets (250k samples) were also integrated into this training phase.
+A summary of the datasets used can be found in Table 5.
+Stage 2: Visual Aids Generation and Understanding We specifically focused on training for visual aids generation.
+Here, "generation" refers to generating visual aids based on a given question, while "understanding" represents its inverse problem.
+We excluded some inverse problems with ambiguous semantics.
+The final total amount of data used is shown in Table 5.
+Training Configuration We used exactly the same hyperparameters in both stages of training.
+We utilized a global batch size of 128 and the AdamW optimizer, configured with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay coefficient of 0.
+The learning rate was set to $2\times 10^{-5}$.
+Both training stages employ a linear warmup for the first 3% of training steps, followed by a cosine decay strategy to a minimum learning rate of 0.
+We simultaneously train both the vision-language connector and the language model.
+The image resolution was set to $336\times 336$, and the visual encoder remained frozen throughout the entire training process.
+In the first phase, we train for 1 epoch on the complete dataset, while in the second phase, we train for 3 epochs.
+The FSD model has approximately 13B trainable parameters and we conducted training using 8 A100 40G GPUs, with Stage 1 requiring approximately 72 hours and Stage 2 requiring 8 hours.
+Appendix C Details of Action Execution
+When utilizing FSD for robotic manipulation tasks, we can select from various visual aids.
+As described in Fig.
+2, these include spatial affordance boxes ($\mathcal{B}$), spatial affordance points ($\mathcal{P}$), and object-centric visual traces ($\boldsymbol{\tau}$).
+The choice of visual aids dictates the subsequent motion planning strategy.
+Motion Planning with Spatial Affordances For spatial affordance boxes ($\mathcal{B}$), the target point for manipulation is derived by sampling the center of the box.
+In the case of spatial affordance points ($\mathcal{P}$), a point is directly sampled.
+When relying on spatial affordance information, whether from boxes or points, we employ CuRobo Sundaralingam et al.
+(2023) as our motion planner.
+CuRobo is responsible for generating collision-free paths that guide the robot’s end-effector to the inferred target affordance point.
+Motion Planning with Object-Centric Visual Traces When leveraging object-centric visual traces ($\boldsymbol{\tau}$), the process involves mapping 2D visual traces into 3D space and then interpolating these discrete points to form a complete motion trajectory in SE(3) space.
+The detailed procedure is as follows:
+We directly acquire 2D keypoint information, denoted as $k_{i}=(u_{i},v_{i})\in\mathbb{R}^{2}$, where $u_{i}$ and $v_{i}$ represent the x and y coordinates of the $i$-th point in the image, for $i\in[1,T]$.
+Initial depth information, $d_{i}\in\mathbb{R}$, is obtained from a depth camera.
+Using the Pinhole camera model Hartley and Zisserman (2004), we can transform these 2D keypoints into 3D Cartesian coordinates $P_{i}=(x_{i},y_{i},z_{i})\in\mathbb{R}^{3}$ via:
+$s_{i}\begin{bmatrix}u_{i}\\ v_{i}\\ 1\end{bmatrix}=\begin{bmatrix}f_{x}&0&c_{x}\\ 0&f_{y}&c_{y}\\ 0&0&1\end{bmatrix}\begin{bmatrix}x_{i}\\ y_{i}\\ z_{i}\end{bmatrix}$
+[…]
+However, a naive use of the raw depth values $d_{i}$ obtained from the depth camera often results in trajectories that closely hug the object’s surface, which is undesirable for robust robot manipulation.
+To address this, we formulate an optimization problem to address it.
+We fix the depth values for the start and end points of the path ($d_{1}$ and $d_{T}$) and optimize the intermediate depth values $d_{2:T-1}$ to minimize the total Euclidean distance between consecutive points in Cartesian space.
+The objective function for this optimization is:
+$\hat{d}_{i}=\arg\min_{d_{2:T-1}}\sum_{i}{d(P_{i},P_{i+1})}$
+where $d(P_{i},P_{i+1})$ represents the Euclidean distance between points $P_{i}$ and $P_{i+1}$.
+We employ a gradient descent method from scipy library to optimize this objective function.
+This refined approach allows for more robust and practical robot motion planning by addressing the limitations of raw depth data and providing a structured framework for integrating various visual aids.
+Appendix D Details of Visual Aids Generation Benchmark
+To address the gap in evaluating visual assistance signal generation, we established VABench, a comprehensive benchmark.
+VABench comprises two distinct tasks: VABench-Point and VABench-VisualTrace, each featuring 300 meticulously hand-annotated questions.
+These tasks require models to infer visual auxiliary information solely from natural language instructions, mimicking everyday human commands.
+For VABench-Point, we provide ground truth bounding boxes for each question.
+The model’s performance is then assessed by calculating the proportion of predicted points that fall within the target region.
+For models that only output bounding boxes, we explored two scoring methods: Intersection Over Union (IOU) and uniformly sampling points within the predicted box.
+We ultimately opted for the latter approach, as the physical interaction between a robotic arm and an object in real-world tasks is fundamentally determined by specific point coordinates.
+For VABench-VisualTrace, we provide ground truth trajectories consisting of eight points.
+When the predicted trajectory length deviates from the ground truth, we employ interpolation to align their lengths.
+To ensure consistent evaluation across varying image resolutions, all coordinates are uniformly normalized to a range of 00 to $1000$.
+Subsequently, we employ a combination of Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and GPT Score to comprehensively evaluate performance.
+Here, we detail the design philosophy and evaluation procedure of the GPT Score.
+Given that multiple valid solutions can exist for each task instruction, relying solely on trajectory similarity to the ground truth as a scoring criterion can be one-sided.
+Therefore, we established detailed evaluation criteria to simulate realistic human assessment.
+Based on these rules, we introduced a visualized trajectory scoring method leveraging Multimodal Large Language Models (MLLMs), termed GPT-score.
+Specifically, we designed an evaluation prompt to guide GPT-4.1 in assessing predicted object manipulation trajectories based on both task instructions and visual inputs.
+The prompt provides clear instructions and criteria, positioning the model as an expert evaluator in robotic manipulation and visual reasoning.
+Each evaluation instance consists of a task instruction and an accompanying image that visualizes the predicted trajectory, where a red circle indicates the start point and a blue diamond marks the end point.
+The model is instructed to assess the trajectory according to three key criteria: (1) task alignment and success, determining whether the predicted path correctly fulfills the instruction by starting and ending in appropriate locations; (2) feasibility, evaluating the physical plausibility and smoothness of the motion; and (3) obstacle avoidance, considering whether the trajectory avoids potential collisions.
+The prompt emphasizes that completing the task correctly is the most important factor; any major deviation in goal achievement results in a low score, even if the trajectory appears smooth or feasible.
+Then, the model returns a structured response consisting of a numerical score from 1 to 10, along with a concise explanation.
+Scores are interpreted based on task success and quality: low scores (1–4) indicate failure to accomplish the task, mid-range scores (6–8) reflect successful but imperfect trajectories, and high scores (9–10) are reserved for trajectories that are both accurate and high-quality.
+This scoring scheme allows for nuanced, human-like evaluation that integrates both semantic understanding and visual reasoning.
+By leveraging this multimodal prompt framework, GPT-score enables a robust and interpretable evaluation process that aligns closely with human judgment, overcoming the limitations of purely geometric or distance-based metrics.
+The complete prompt is presented as follows:
+Complete Prompt for GPT Score Metrics You are an expert evaluator in robotic manipulation and visual reasoning.
+Your job is to assess the quality of predicted trajectories based on task instructions and visual inputs.
+You are given: - A task instruction describing an object manipulation task.
+- An image showing a predicted trajectory.
+**Note:** - In the image, the red circle indicates the start point, and the blue diamond indicates the end point.
+- The trajectory represents the predicted movement path of the manipulated object, not the robot or end-effector.
+- You should **evaluate the predicted trajectory as a proposed motion for the object that is supposed to be moved**, based on the task instruction — **not based on the static positions of objects in the image**.
+The objects have not actually moved.
+**Evaluation Criteria (listed in order of importance):** 1.
+**Task Alignment and Success (most important)** - Does the trajectory clearly and accurately fulfill the task instruction?
+- **The trajectory must start at the correct location and end at a target position that aligns with the task goal.** - Large deviations in the starting or ending point (e.g., wrong object, wrong destination, or stopping short of the goal) should result in a low score, even if the rest of the trajectory is smooth.
+- If the task is not accomplished (due to incorrect goal interpretation or spatial execution), the score should be low regardless of other qualities.
+2.
+**Feasibility** - Is the movement physically plausible, smooth, and continuous?
+- Are there any unrealistic discontinuities, sharp turns, or impossible transitions?
+- Even if the movement is feasible, it should not receive a high score if the task is not completed.
+3.
+**Obstacle Avoidance / Safety** - Does the trajectory reasonably avoid collisions with surrounding objects?
+- Minor risks may be tolerated if the task is completed successfully, but major or clear collisions should reduce the score.
+**Scoring Guideline:** - If the task is **not accomplished**, or if the start or end point is significantly incorrect, the score should typically be **4 or below**.
+- If the task is completed but the trajectory has issues (e.g., roughness, minor risk of collision), a score in the **6–8** range is appropriate.
+- A **score of 9–10** should be given only when the trajectory clearly completes the task, with good start/end accuracy, smooth motion, and reasonable safety.
+Based on these criteria, provide a single overall score from 1 (very poor) to 10 (excellent), reflecting how well the task is accomplished.
+Respond strictly in the following format: Score: <1-10> Explanation: <brief justification> The task instruction is: {task_instruction} Please give your response.
+Table 5: Details of the training data for FSD
+Stage Task Datasets Samples Data Sources
+Stage 1 GeneralQA (Caption, VQA, OCR, RegionVQA, Conversation, Grounding, Text) ShareGPT4V, VQAv2, OCR-VQA, Visual7W, ST-VQA, RefCOCO/+/g, VG, AS-Core, AS-V2, TextVQA, Visual7W 838k LLaVA Liu et al.
+(2024a), ASMv2 Wang et al.
+(2025)
+General Spatial Reasoning KITTI, 2D3DS, ObjectRef, RegionRef, VSR, CLEVR, CLEVR-Math, SUPER-CLEVR, RAVEN 295k SpatialBot Cai et al.
+(2024), RoboPoint Yuan et al.
+(2024b), SAT Ray et al.
+(2024), LLaVA-OneVision Li et al.
+(2024a)
+Embodied Spatial Reasoning FSD Level 1 (145k), Level 2 (86k) and Level 3 (19k) 250k FSD
+Stage 2 Spatial Affordance Generation & Understanding FSD Level 4 (24k) 24k FSD
+Visual Trace Generation & Understanding FSD Level 5 (26k) 26k FSD
+Appendix E Details of Benchmarks and Baselines
+For general spatial reasoning tasks, the answers are typically multiple-choice questions with clear options.
+However, some spatial reasoning models show reduced instruction-following ability after fine-tuning, preventing them from directly outputting the correct option.
+To address this, we use a lenient matching rule, considering an answer correct if it includes either the correct content or the corresponding option.
+For object/region reference tasks, we carefully fine-tuned and used a tailored prompt for each model.
+Most models, such as GPT-4o and ASMV2 Wang et al.
+(2025), cannot directly output specific points.
+Similar to the validation process for RoboPoint Yuan et al.
+(2024b), we also found that using in-context learning to specify point output formats resulted in worse performance compared to directly outputting bounding boxes.
+Therefore, for these models, we instructed them to output bounding boxes directly.
+From these bounding boxes, we either uniformly sampled nine points or took the midpoint.
+Then, we calculated the proportion of points within the specified region to determine the final average accuracy.
+For the VABench-VisualTrace task, due to a lack of strong baselines, we developed an additional Transformer-based prediction model, trained from scratch using the same data, which we named DINOv2 Predictor.
+In the DINOv2 Predictor, the visual encoder uses a pre-trained DINOv2 Oquab et al.
+(2023), encoding images to output a $(196,768)$ feature vector.
+The language encoder uses a pre-trained T5-Base Raffel et al.
+(2020), outputting $(32,768)$.
+These are concatenated with learnable embeddings $(8,768)$ and passed through a Transformer encoder base.
+The output of the learnable embeddings is then read and passed through a linear layer to predict eight points.
+During training, we keep the language encoder fully frozen and train the visual encoder along with the other remaining parameters.
+Appendix F More Visualizations and Examples
+We present the prediction results of FSD on Where2place Yuan et al. (2024b), Roborefit Lu et al. (2023), and VABench in Fig. 12, Fig. 13, and Fig. 14, respectively.
+Figure 12: Visualization of visual aids generated by FSD in the Where2Place benchmark.
+Figure 13: Visualization of visual aids generated by FSD in the RoboRefit benchmark.
+Figure 14: Visualization of visual aids generated by FSD in the VABench benchmark.
+FSD can generate three types of visual aids based on task instructions for novel tasks and scenarios.
+1st-2nd row: visual trace; 3rd-4th row: affordance points; 5th-6th rows: affordance bounding box.
+Appendix G Real World Experiment Results
+In our real-world desktop manipulation tasks, we used an xArm 6 robotic arm for evaluation.
+This setup included an Intel RealSense L515 LiDAR camera and a force-torque sensor on the xArm to enable compliance control, which improved interaction with the environment.
+A computer running Ubuntu 24.04 and equipped with an NVIDIA GTX 1660 was directly connected to the arm and camera to execute low-level control policies.
+Notably, a single RealSense L515 depth camera was sufficient for task completion, especially when performing visual trace execution.
+This approach eliminated the need for object segmentation and 3D mapping; instead, we directly mapped 2D visual trajectories to 3D for execution, with no strict requirements on depth information accuracy.
+Demonstrations are available in Fig.
+6 and on our website.
+Appendix H Prompt for Using FSD Model
+Generate Spatial Affordance Points & Bounding Box You are currently a robot performing robotic manipulation tasks.
+Your task instruction: {Task Instruction} .
+Observe the image, use 2D points and bounding box to mark the target location where the manipulated object will be moved.
+In your answer, use <box>[[x1, y1, x2, y2]]</box> to present the bounding box of the target region, and use <point>[[x1, y1], [x2, y2], …]</point> to mark the points of the free space.
+Generate Visual Trace You are currently a robot performing robotic manipulation tasks.
+Your task instruction: {Task Instruction}.
+Observe the image, use 2D points to mark the manipulated object-centric waypoints to guide the robot to manipulate the object.Typically, the waypoints consists of an ordered sequence of eight 2D points.
+The format is <point>[[x1, y1], [x2, y2], …]</point>.
+Appendix I Comparison of FSD and RoboBrain
+Both FSD and RoboBrain Ji et al.
+(2025) have the capability to generate visual trace.
+However, RoboBrain tends to produce agent-centric visual trace, whereas FSD generates object-centric visual trace.
+FSD adopts a task-centric design principle, allowing it to perform effectively even in more heterogeneous ontological scenarios, including those that completely lack robotic arms in the image, thus exhibiting stronger generalizability.
+Due to the differences in the methods of generating visual trace, we conducted several sets of visual trajectory visualizations for qualitative analysis, as shown in Fig.
+15.
+Under the same zero-shot setting, the visual trace generated by FSD has higher accuracy compared to those generated by RoboBrain, confirming that FSD’s reasoning-based pipeline possesses greater generalizability when facing unknown tasks.
+Figure 15: Comparison of generated visual traces between FSD and RoboBrain.
+Appendix J Future Works
+We have made preliminary attempts to use visual aids as intermediate states in FSD, achieving promising results in object/target/region reference and actual manipulation task execution.
+Future work can focus on the following aspects to further enhance the applicability of this paradigm: 1.
+Task Decomposition for Complex and Long-Horizon Instructions: The current version of FSD primarily targets clear and explicit language instructions.
+When dealing with long-horizon tasks or ambiguous/complex instructions, the model needs to decompose them into atomic, executable sub-tasks.
+We believe that decomposing instructions into a sequence of visual aids to guide each sub-task execution is a promising avenue.
+2.Downstream Execution and Visual-Aid-Guided Control: Currently, FSD relies mainly on training-free motion planning methods for downstream execution.
+In extremely complex or dynamic scenarios, this may lead to a bottleneck in success rates.
+A potential improvement is to use the generated visual aids as explicit guidance for downstream VLA models, replacing language-conditioned training.
+Several preliminary studies Zheng et al.
+(2024), Bharadhwaj et al.
+(2024), Li et al.
+(2025) have shown that, for robotic manipulation tasks, affordance and visual trajectories provide more effective guidance than language prompts.
+3.
+Extending from 2D to 3D Visual Aids: At present, FSD focuses on predicting 2D visual aids, similar to the representation used in Referring Expression Comprehension (REC) and related tasks, which leverages the general visual understanding and reasoning capabilities of VLMs.
+However, as task and scene complexity increases, predicting 3D visual traces may prove to be a more effective solution, and we identify this as an important direction for future research.